ThisGirlPretty:

Robin can mow a lawn in 3 hours, while Brady can mow the same lawn in 4 hours. How many hours would it take for them to mow the lawn together? seven twelfths twelve sevenths 7 12

ThisGirlPretty:

Okie :D

ThisGirlPretty:

Huh? @Hero

Hero:

There is a formula you can use to get the right answer to this question.

Hero:

Let M = the time it takes the first person to mow the lawn alone. Let N = the time it takes the second person to mow the lawn alone. Then \(\dfrac{M \times N}{M + N}\) represents the time it takes if both M and N mowed the lawn together.

ThisGirlPretty:

Ok

Hero:

You do get the same answer @Ultrilliam posted but it's better if you know the formula that way you know how to solve these on your own.

ThisGirlPretty:

That's true because I have a dba coming up lol

Hero:

Also, it's better if you understand what the answer means in terms of hours. Twelve sevenths is equivalent to \(1 + \dfrac{5}{7}\)

Hero:

In other words, if Robin and Brady mowed the lawn together, they would finish in one and five-seven hours or approximately two hours.

ThisGirlPretty:

Ok I see now, thank you @Hero